# Exploring Training and Inference Scaling Laws in Generative Retrieval

### v1

@@ 7. Conclusion and Future Work
In this paper, we explored the scaling laws of generative retrieval, analyzing how model size, training data size, model architecture, and retrieval methods impact retrieval performance. Using Contrastive Generation Loss as our primary metric, we found a clear power-law relationship between model size and retrieval performance for n-gram-based methods. Additionally, we observed consistent scaling trends with increasing data size for both n-gram-based and codebook-based methods. A comparison between architectures showed that LLaMA consistently outperformed T5 under identical experimental conditions, highlighting LLaMA’s superior capability for advancing generative retrieval.
Beyond model and data scaling, we also investigated inference scaling, specifically for n-gram-based methods. Our results revealed that increasing inference computation followed a distinct power-law trend, significantly boosting retrieval performance. This unique inference scaling offers complementary opportunities for optimizing retrieval efficiency and performance. Collectively, our findings validate the effectiveness of systematically scaling model capacity, training data, and inference computation, suggesting promising pathways to further enhance generative retrieval methods.
While our research provides a foundation for understanding the scaling laws of generative retrieval, several directions for future exploration remain. First, further investigation into scaling laws for codebook-based methods is warranted, particularly in the context of model size. Exploring advanced training objectives could better capture the potential scaling benefits of these methods, which were not fully realized under our experimental settings. Second, with access to greater computational resources, revisiting extensive datasets could enable deeper investigations into the scaling laws of generative retrieval, particularly for methods that require larger-scale training. Finally, while this study focused on n-gram-based methods for inference scaling, exploring similar scaling strategies for codebook-based methods, such as increasing beam size, remains an intriguing direction for future research.

### v2

@@ 7. Conclusion and Future Work
In this paper, we explored the scaling laws of generative retrieval across training and inference dimensions. We analyzed how model size, training data size, and retrieval methods influence performance. Using Contrastive Generation Loss as the metric, we found a clear power-law relationship between model size and retrieval performance for n-gram-based methods. Additionally, we observed consistent scaling trends with increasing data size for both n-gram-based and codebook-based methods. A comparison between architectures showed that LLaMA consistently outperformed T5 under identical experimental conditions, highlighting LLaMA’s superior capability for advancing generative retrieval. Beyond training scaling, we also investigated inference scaling for n-gram methods and found that increasing inference-time computation followed a power-law trend, offering a complementary axis for performance gains. Collectively, our findings validate the effectiveness of systematically scaling model capacity, training data, and inference computation, suggesting pathways to further enhance generative retrieval methods.
While our study offers a foundation for understanding the scaling laws of generative retrieval, several directions remain for future exploration. First, adopting more effective training objectives, such as learning-to-rank losses, may improve retrieval performance by better aligning model optimization with the retrieval task. Second, for codebook-based methods, the limited model-size scaling observed in our experiments suggests a need for larger training datasets and longer training schedules to fully realize their potential. Finally, although our inference scaling analysis focused on n-gram-based methods, codebook-based methods may also benefit from increased inference-time computation, such as using larger beam sizes—offering a valuable direction for extending the scope of inference scaling in generative retrieval.
Acknowledgements.
This research is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-2024-002), Research Grants Council of Hong Kong(PolyU/15209724) and PolyU internal grants (BDWP). Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore.